Circular No. 7994

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002gn, go, gp, gq, gr, gs, gt, gu, 2002gv G. Miknaitis, University of Washington, reports the detection by the

Sloan Digital Sky Survey (SDSS) of nine supernovae in images taken on Oct. 10.4 UT with the SDSS 2.5-m telescope at Apache Point Observatory (APO). The new objects were detected by subtraction against images taken on 2001 Nov. 12 (limiting mag r = 22.5). SDSS r magnitudes and each object’s offset from the center of its host galaxy (SN 2002gq has no apparent host) are given below. CCD spectra taken by B. Leibundgut and J. Sollerman at the Very Large Telescope on Oct. 11 (first five) and 12 (SNe 2002gs, 2002gt) reveal the following supernova types, approximate time before (−) or after (+) maximum light, and redshifts: SN 2002gn, Ia, +13 days, 0.044 (from Mgb and Na I D lines in the host spectrum); SN 2002go, Ia, −2 days, 0.236; SN 2002gp, Ia, +9 days, 0.18; SN 2002gq, peculiar II, 0.20 (wide emissions of Hα and Hβ and no P-Cyg lines); SN 2002gr, Ia, +2 weeks, 0.09 (from galaxy emission lines); SN 2002gs, Ia, +2 weeks, 0.33; SN 2002gt, Ia, +10 days, 0.37. SNe 2002gu and 2002gv were confirmed by CCD imaging carried out by Miknaitis and A. Miceli at the APO 3.5-m telescope on Oct. 12.

SN  2002 UT  α_{2000}  δ_{2000}  r  Offset
2002gn Oct. 10.4 1^{h}56^{m}54.^{s}56 − 1^{o}06′49′′.0 19.8 1′′.8 E, 4′′.0 N
2002go  10.4 2 45 43.16 − 0 37 44.7 21.3 3′′.7 E, 2′′.7 N
2002gp  10.4 1 52 36.76 − 0 13 31.5 21.2 2′′.4 W, 9′′.5 N
2002gq 10.4 2 10 29.69 + 0 32 16.9 21.0 —
2002gr  10.4 2 27 13.22 + 0 53 57.8 21.8 3′′.3 E, 1′′.8 N
2002gs  10.4 1 39 56.44 − 0 45 01.2 21.4 1′′.0 W, 1′′.0 N
2002gt  10.4 2 40 29.78 − 0 22 46.0 21.7 2′′.0 W, 2′′.0 S
2002gu  10.4 2 31 48.56 − 1 09 19.7 20.9 4′′.0 E, 6′′.5 N
2002gv  10.4 1 52 17.48 + 0 12 17.3 20.6 2′′.6 W, 2′′.0 S

SUPERNOVA 2001ig IN NGC 7424 P. Chandra and A. Ray, Tata Institute of Fundamental Research,

Mumbai; and S. Bhatnagar, National Radio Astronomy Observatory, Socorro, report the detection of SN 2001ig at 1396.75 MHz (flux density 11.7 ± 1.5 mJy) with the Giant Meterwave Radio Telescope (GMRT) on Sept. 25.8 UT. SN 2001ig is 21′′.3 away from a radio source of flux density 38.4 ± 1.6 mJy. In an earlier GMRT observation on 2001 Dec. 31, the 2σ upper limit of flux density of the supernova in the 610-MHz band was 2 mJy.

2002 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green